Determine the equation of g(x) that results from translating the function
f(x)=(x+8)^2 to the right 14 units.
A. g(x)=(x-6)^2
B. g(x)=(x+22)^2
C. g(x)=(x+8)^2-14
D. g(x)=(x+8)^2+14

Answer by Edwin McCravy(20055)

f(x) = (x+8)2

Replace x by x-14

g(x) = f(x-14) = (x-14+8)2 = (x-6)2

Edwin